OpenStudy (anonymous):

Solve the equation. Square root of 5x-1 = Square root of 4x+9

OpenStudy (anonymous):

\[\sqrt{5x-1}=\sqrt{4x+9}\]

OpenStudy (the_fizicx99):

Start by doing the inverse, square both sides to cancel out the square root. \(\ (\sqrt{5x - 1})^2 = (\sqrt{4x + 9})^2\) 5x - 1 = 4x + 9 +1 +1 5x = 4x + 10 -4x -4x x = 10 -------------- \(\\ \sqrt{5(10) - 1} = \sqrt{50 - 1} = \sqrt{49} = 7 \) \(\ \sqrt{4(10) + 9} = \sqrt{40 + 9} = \sqrt{49} = 7 \) They both equal 7 when you check if 10 is correct. Therefore x does indeed equal 10 \( \color{green}\checkmark\)

OpenStudy (anonymous):

Thanks! Solve the equation for x. Identify any extraneous solutions \[y=\sqrt{-6y}\]

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

x? There's only a y

OpenStudy (anonymous):

Sorry Solve the equation for y. Identify any extraneous solutions

OpenStudy (anonymous):

@Hero

OpenStudy (the_fizicx99):

Are you sure it's y = √-6y ?

OpenStudy (anonymous):

positive the answer choices are stuff like 1.0 and -6 are solutions to the original equation

OpenStudy (the_fizicx99):

Really? It's a constant back and fourth loop, y = \(\ \sqrt{-6y} \) square both sides and divide by -6, you would just get; y^2/-6 = y And if you solved for the left side you would end up with the original equation. If anything y = 0 0 = √-6(0) 0 = √0 = 0 .-.

OpenStudy (the_fizicx99):

Waait, I \(\ think\) its -6 ♦o♦

hero (hero):

@tHe_FiZiCx99, how come you didn't square both sides like you did with the first problem?

OpenStudy (the_fizicx99):

What do you mean, I did?

hero (hero):

lol, you're right

hero (hero):

You did something wrong though

OpenStudy (the_fizicx99):

I'm brain storming these in my head, and writing them along. I might have some typos.

hero (hero):

You should have ended up with y^2 + 6y = 0 y(y + 6) = 0 y = 0 y = -6

OpenStudy (the_fizicx99):

yeah, typo >_<

OpenStudy (the_fizicx99):

I actually did it another way in my head :/

OpenStudy (the_fizicx99):

Yeah, y = √6 * √-y = y^2 = 6 * (-y) y^2 = 6*(-y) /-y /-y -y = 6 /-1 /-1 y = -6 y = 0 Yeah somehow I managed to do this in my head and forget it for a bit :<

hero (hero):

The only problem is, you're not allowed to divide by the variable you are solving for. The reason is because it actually eliminates one of the solutions.

OpenStudy (the_fizicx99):

Yeah, I thought about the zero's but technically I assumed there's an understood 0

OpenStudy (the_fizicx99):

Fun problem nonetheless :>

hero (hero):

I understand what you're trying to say If y = 0 then \(0 = \sqrt{-6(0)}\) 0 = 0 However, specificially with regard to solving for y, there's a proper solving methodology and mathematicians have agreed that dividing by the variable you're solving for is not proper for two reasons: 1. You eliminate a solution 2. y = 0 and dividing by zero is frowned upon.

hero (hero):

There's an understood zero, but that doesn't justify dividing by zero

OpenStudy (the_fizicx99):

Oh that part, yeah :/ you can't divide by 0's

OpenStudy (the_fizicx99):

Thanks for clarifying that though :P

hero (hero):

yw

OpenStudy (anonymous):

...............mind blown haha anyways i dont get what the answer would be?

OpenStudy (anonymous):

y=0 & y=-6

hero (hero):

Actually, y = 0 may be the only correct solution.

hero (hero):

If we check with y = -6 we get the following: \(-6 = \sqrt{-6(-6)}\) \(-6 = \sqrt{36}\) \(-6 = 6\) FALSE

hero (hero):

So y = 0 is therefore the ONLY correct solution.

OpenStudy (the_fizicx99):

"Sorry Solve the equation for y. Identify any extraneous solutions" y = -6 Would be considered an \(\ extraneous \) solution

OpenStudy (anonymous):

alright thank you so much could you help with just one more please?

OpenStudy (the_fizicx99):

Sure, I'll try

OpenStudy (anonymous):

\[\sqrt{x=-3}\]

OpenStudy (anonymous):

Solve the equation for x. Identify any extraneous solutions

OpenStudy (the_fizicx99):

o_o that's it?

hero (hero):

lol, really?

OpenStudy (the_fizicx99):

I think you missed something >.>

hero (hero):

Maybe he means \(\sqrt{x} = \sqrt{-3}\)

hero (hero):

In that case, you start by squaring both sides.

hero (hero):

You might run into a problem while checking though

OpenStudy (anonymous):

sorry \[\sqrt{x}=-3\]

hero (hero):

Either way, still square both sides and run into a problem while checking

OpenStudy (the_fizicx99):

Square both sides, x = 9 >.> √9 = 3 not -3, that's extraneous

OpenStudy (the_fizicx99):

I'm curious ot know the +/- sign with alt keys :3

OpenStudy (the_fizicx99):

to*

OpenStudy (anonymous):

so their is no real solution

OpenStudy (the_fizicx99):

Yeah, there's no solution, "That's extraneous"